OpenStudy (flowey15):

Rearrange the formula P = 2(L + b) for b. A) b = (P - 2l) B) b =p/2-2l C) b=P/2+2L D) b=1/2(p-2l)

OpenStudy (onyxxelite):

what do you think it is

OpenStudy (onyxxelite):

i am not allowed to give direct answers

OpenStudy (flowey15):

i didnt get one of those, i got b=p/2-L

OpenStudy (onyxxelite):

ok let me check solve real quick

OpenStudy (flowey15):

but none ae the answer i got

OpenStudy (onyxxelite):

ok then we must have messed up somewhere explain your steps to me

OpenStudy (flowey15):

divide by 2 on both sides= 9/2=L+b subtract L= p/2-L

OpenStudy (flowey15):

b= p/2-L

OpenStudy (onyxxelite):

Well, i am going to assume that the correct answer is B. only because it is the closest match to yours.

OpenStudy (flowey15):

WHAT DID YOU GET WHEN YOU WORKED IT OUT?

OpenStudy (onyxxelite):

im not allowed to give direct answers but i "THINK" it's B